# Trivial Isomorphisms between Reduced Products

### v2

@@ 6.2. Continuum Hypothesis and the strong failure of rigidity
(2) Let $\mathcal{L}_{n}$ and $\mathcal{L}^{\prime}_{n}$ be sequences of linear orderings that satisfy $\limsup_{n}|\mathcal{L}_{n}|=\infty$. As in (1a) and (1b), it suffices to prove that $\prod_{n}\mathcal{L}_{n}/\operatorname{Fin}$ and $\prod_{n}\mathcal{L}^{\prime}_{n}/\operatorname{Fin}$ are elementarily equivalent. By the Feferman–Vaught theorem it suffices to prove that the theories of every sequence of linear orderings the supremum of whose sizes is $\aleph_{0}$ converge in the logic topology (this implies that they all necessarily converge to the same limit).
@@ 7. Rigidity of products
If ${\mathcal{G}}_{n}$, ${\mathcal{H}}_{n}$, for $n\in{\mathbb{N}}$, are sequences of random graphs and $\Phi\colon\prod_{n}{\mathcal{G}}_{n}\to{\mathcal{H}}_{n}$ is an isomorphism, then there are a permutation $\pi$ of ${\mathbb{N}}$ and isomorphisms $\sigma_{n}\colon{\mathcal{G}}_{n}\to{\mathcal{H}}_{\pi(n)}$ for $n\in{\mathbb{N}}$ such that, for all $x\in\prod_{n}\mathcal{G}_{n}$, $\Phi(x)(\pi(n))=\sigma_{n}(x(n))$.
@@ 8. Concluding remarks
First, what other theories recognize coordinates in reduced products? Can recognizing coordinates be re-framed in terms of the existence of a copy of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ in the $\operatorname{eq}$ (see [23, p. 151]) of the reduced product at hand? Or, are theories that recognize coordinates characterized by the existence of local copies of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$, like the ones obtained for ramified sets and random graphs (§2.2 and §2.3), in reduced products of its models? Long story short, is there a natural dividing line for theories $T$ such that the reduced products of models of $T$ recognize coordinates (compare with the order property/stability as in [32])? An easy observation is that if $T$ is a recognizes coordinates and $T’$ ’ is a theory in an expanded language extending $T$ then $T’$ ’ recognizes coordinates as well.
@@ Appendix A From $\sigma$-Borel to continuous
Suppose that $\Phi\colon{\mathcal{P}}({\mathbb{N}})\to{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ is a homomorphism, ${\mathbb{N}}=\bigsqcup_{n}A_{n}$, and there are Borel-measurable functions $F_{n}\colon{\mathcal{P}}({\mathbb{N}})\to{\mathcal{P}}({\mathbb{N}})$, for $n\in{\mathbb{N}}$, whose graphs cover the graph of a lifting of $\Phi$. Then for all but finitely many $n$ the restriction of $\Phi$ to $A_{n}$ has a continuous lifting.

### v3

@@ 6.2. Continuum Hypothesis and the strong failure of rigidity
(2) Let $\mathcal{L}_{n}$ and $\mathcal{L}^{\prime}_{n}$ be sequences of linear orderings with endpoints that satisfy $\limsup_{n}|\mathcal{L}_{n}|=\limsup_{n}|\mathcal{L}^{\prime}_{n}|=\aleph_{0}$. As in (1a) and (1b), it suffices to prove that $\prod_{n}\mathcal{L}_{n}/\operatorname{Fin}$ and $\prod_{n}\mathcal{L}^{\prime}_{n}/\operatorname{Fin}$ are elementarily equivalent. By the Feferman–Vaught theorem it suffices to prove that the theories of every sequence of linear orderings the supremum of whose sizes is $\aleph_{0}$ converge in the logic topology (this implies that they all necessarily converge to the same limit).
@@ 7. Rigidity of products
If ${\mathcal{G}}_{n}$, ${\mathcal{H}}_{n}$, for $n\in{\mathbb{N}}$, are sequences of random graphs and $\Phi\colon\prod_{n}{\mathcal{G}}_{n}\to\prod_{n}{\mathcal{H}}_{n}$ is an isomorphism, then there are a permutation $\pi$ of ${\mathbb{N}}$ and isomorphisms $\sigma_{n}\colon{\mathcal{G}}_{n}\to{\mathcal{H}}_{\pi(n)}$ for $n\in{\mathbb{N}}$ such that, for all $x\in\prod_{n}\mathcal{G}_{n}$, $\Phi(x)(\pi(n))=\sigma_{n}(x(n))$.
@@ 8. Concluding remarks
First, what other theories recognize coordinates in reduced products? Can recognizing coordinates be re-framed in terms of the existence of a copy of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ in the $\operatorname{eq}$ (see [23, p. 151]) of the reduced product at hand? Or, are theories that recognize coordinates characterized by the existence of local copies of ${\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$, like the ones obtained for ramified sets and random graphs (§2.2 and §2.3), in reduced products of its models? Long story short, is there a natural dividing line for theories $T$ such that the reduced products of models of $T$ recognize coordinates (compare with the order property/stability as in [32])? An easy observation is that if $T$ recognizes coordinates and $T’$ ’ is a theory in an expanded language extending $T$, then $T’$ ’ recognizes coordinates as well.
@@ Appendix A From $\sigma$-Borel to continuous
Suppose that $\Phi\colon{\mathcal{P}}({\mathbb{N}})\to{\mathcal{P}}({\mathbb{N}})/\operatorname{Fin}$ is a homomorphism, ${\mathbb{N}}=\bigsqcup_{n}A_{n}$, and there are Borel-measurable functions $F_{n}\colon{\mathcal{P}}({\mathbb{N}})\to{\mathcal{P}}({\mathbb{N}})$, for $n\in{\mathbb{N}}$, whose graphs cover the graph of a lifting of $\Phi$. Then for all but finitely many $n$ the restriction of $\Phi$ to ${\mathcal{P}}(A_{n})$ has a continuous lifting.